OpenStudy (anonymous):

Please Help..... Solve.... what does x equal? 4x^2-11x=0

OpenStudy (anonymous):

\[x(4x-11)=0\] is the first step

OpenStudy (anonymous):

Thank you what do we do next? @satellite73

OpenStudy (goformit100):

it was an easy question.

OpenStudy (anonymous):

is that the answer? x(4x-11)=0 @goformit100

OpenStudy (goformit100):

yes it is x(4x-11)=0

OpenStudy (anonymous):

I forgot to mention were suppose to solve for x... Would that still be the answer? @goformit100

OpenStudy (goformit100):

ok now you can solve it

OpenStudy (anonymous):

How? @goformit100

OpenStudy (goformit100):

equate for 0 now

OpenStudy (anonymous):

I still dont know how? @goformit100

OpenStudy (goformit100):